OpenStudy (anonymous):

150 Grams+ how many Milligrams?

OpenStudy (anonymous):

its supposed to be = how many milligrams

OpenStudy (anonymous):

1000 mg = 1 gram so now multiply both sides by 150 150*1000 mg = 150 *1 grams 150000 mg = 150 grams

OpenStudy (anonymous):

i got 4500? is that correct?

OpenStudy (anonymous):

how did you get 4500?

OpenStudy (anonymous):

victoria11 <<< she got that answer..

OpenStudy (anonymous):

well i got 150,000 mg = 150 grams remember 1000 mg = 1 gram so 150 grams = 1000 * 150 mg

OpenStudy (anonymous):

thats sounds right im gonna put that in

OpenStudy (anonymous):

:) is this online?

OpenStudy (anonymous):

its like, if 1 dozen apples = 12 apples. then 2 dozen = 24 apples. (we multiply both sides by 2) or more generally. n dozen apples = 12*n apples

OpenStudy (anonymous):

okay what about this ? 4 kl = _______ L

OpenStudy (anonymous):

the k in front of L means multiply by 1000 so 4kL means 4 *1000 Liters

OpenStudy (anonymous):

just like when i say 4 dozen apples, the dozen means multiply by 12. so 4 dozen apples means 4*12 = 48 apples. k is short for kilo by the way

OpenStudy (anonymous):

so the answer is 48?

OpenStudy (anonymous):

that was for my example. for your example 4kL = 4000 L , k is short for kilo which means multiply by 1000. L is short for Liters , which is a type of volume like quarts

OpenStudy (anonymous):

so whats the answer?

OpenStudy (anonymous):

4000 L

OpenStudy (anonymous):

so whats the answer to this one 63Kg=_______mg

OpenStudy (anonymous):

when you see K, insert x1000. so 63 Kg = 63 *1000 g = 63000 g. but we want mg. now m in front of g stands for , x .001. so 63 kg = 63,000 g = 63000 (1000 *.001) g = (63,000 * 1000) mg notice that 1000*.001 = 1 so i didnt change the equality

OpenStudy (anonymous):

so 63 kg = 63,000,000 mg , in case you didn't get that

OpenStudy (anonymous):

\[\ge \div \times +\neq \neq \subseteq\]

OpenStudy (anonymous):

srry wrong chat